OpenStudy (skullcandy229):

When the following expression is written in simplest form, what is the coefficient of the variable term? -2.5y + 8y - 10.5y

OpenStudy (anonymous):

y

OpenStudy (anonymous):

well i plugged it into my calculator and got 0 soo

OpenStudy (anonymous):

its not 0.

OpenStudy (anonymous):

-2.5-10.5 is -13

OpenStudy (anonymous):

-13+8=-5

OpenStudy (anonymous):

the answer is -5y

OpenStudy (jango_in_dtown):

-2.5y + 8y - 10.5y =(-2.5+8-10.5)y=-5y

OpenStudy (jango_in_dtown):

again the application of the distributive property

OpenStudy (jango_in_dtown):

so the coefficient is -5

OpenStudy (skullcandy229):

one of these are correct -5 5.5 0 -21

OpenStudy (jango_in_dtown):

@skullcandy229